# Algorithmic Fairness amid Social Determinants: Reflection, Characterization, and Approach

### v1

@@ Abstract
Social determinants are variables that, while not directly pertaining to any specific individual, capture key aspects of contexts and environments that have direct causal influences on certain attributes of an individual. Previous algorithmic fairness literature has primarily focused on sensitive attributes, often overlooking the role of social determinants. Our paper addresses this gap by introducing formal and quantitative rigor into a space that has been shaped largely by qualitative proposals regarding the use of social determinants. To demonstrate theoretical perspectives and practical applicability, we examine a concrete setting of college admissions, using region as a proxy for social determinants. Our approach leverages a region-based analysis with Gamma distribution parameterization to model how social determinants impact individual outcomes. Despite its simplicity, our method quantitatively recovers findings that resonate with nuanced insights in previous qualitative debates, that are often missed by existing algorithmic fairness approaches. Our findings suggest that mitigation strategies centering solely around sensitive attributes may introduce new structural injustice when addressing existing discrimination. Considering both sensitive attributes and social determinants facilitates a more comprehensive explication of benefits and burdens experienced by individuals from diverse demographic backgrounds as well as contextual environments, which is essential for understanding and achieving fairness effectively and transparently.
@@ 1 Introduction
Structural injustice refers to circumstances in which social practices, social structures, or the environment reinforce and compound prior histories of injustice [20, 114, 137, 128, 107, 6]. We use the term “social determinants” to refer to the specific aspects of social practices, social structures, or the environment that have a profound impact on individuals’ opportunities, behaviors, and outcomes. When members of specific demographic groups have been the subject of histories of unjust treatment, their demographic membership often correlates with circumstances in which they face significant social impediments [47, 134, 104, 135, 27]. Because social determinants are features of places, institutions, policies, or practices, they persist even if animuses that cause unjust treatment have been subject to significant reform. Their effects may not be tied directly to demographic group membership but to broader traits (such as income level or job type) or to geographic areas. As a result, individuals within the same demographic group, depending on their unique circumstances, may experience different levels of (dis)advantage due to intersecting social determinants, e.g., various environmental impacts on health in different geographic locations [33, 136, 123]. Conversely, individuals from different demographic groups in the same geographic neighborhood may encounter similar impediments, e.g., poverty and pollution in the neighborhood, lack of educational resource in the community [34, 127, 106].
Previous research on algorithmic fairness has focused on sensitive attributes, e.g., race, sex, gender, and age [105, 75, 35, 84, 89, 130, 21, 30, 77, 82, 145, 94, 125]. Various fairness metrics that are directly defined upon sensitive attributes are proposed to estimate or bound empirical violations of fairness, based on observational statistics [17, 52, 140], causal properties and/or quantities [67, 72, 86, 28, 36], and dynamic modelings [74, 146, 58, 124]. In terms of auditing potential fairness violations, the focus on sensitive attributes is natural since these are the features in virtue of which individuals might be subject directly to unfair treatment or might experience disproportionate burdens. However, the goal of mitigation goes beyond auditing fairness violations by seeking to intervene in ways that will reduce burdens and promote fairer outcomes in the future. Both sensitive attributes and social determinants play important roles in the underlying causal mechanism, and therefore, need to be explicitly addressed when designing and evaluating mitigation strategies.
Our contributions can be summarized as follows:
  • We identify a critical gap in algorithmic fairness by drawing on cross-disciplinary perspectives on social determinants. We find that existing approaches primarily focus on sensitive attributes, both in technical methods and data processing practices, largely overlooking social determinants.
  • Through a concrete setting of college admissions, we demonstrate how incorporating social determinants, even with geographic region as a simple surrogate, offers quantitative leverage and recovers nuanced insights that resonate with prior qualitative debates in jurisprudence.
  • We provide empirical evidence of real-world disparities in the interplay between sensitive attributes and social determinants. We also demonstrate how to link existing datasets to socioeconomic status indicators using address information, facilitating the development of richer fairness benchmarks.
@@ 2 Cross-Disciplinary Engagement with Social Determinants
In this section, we review and reflect on the cross-disciplinary engagements with social determinants.111 Due to space limit, we provide further discussions on related works in Appendix A. In Section 2.1, we clarify definitions of sensitive attributes and social determinants. In Section 2.2, we summarize discussions on social determinants from the literature of political philosophy, economics, sociology, and healthcare. In Section 2.3, we review approaches and data processing practices in the algorithmic fairness literature, with special attention to sensitive attributes and social determinants.
2.1 Definitions: Sensitive Attributes and Social Determinants
@@
A social determinant is a variable representing an aspect of the data generating process, in which one or more characteristics of the context (e.g., social practices, social structures, and environments) where individuals live or operate, that are not an attribute of any specific individual, have direct influence on individual’s attributes.
Sensitive attributes are stable identifiers arising from longstanding legal and moral frameworks, that can be uniquely ascribed to an individual [105, 75, 35, 84, 89, 130, 21, 30, 77, 57, 82, 145, 94, 125, 109]. Examples include sex, race, ethnic group, disability status, religion, and so on. Social determinants refer to external conditions that influence an individual’s opportunities, behaviors, and outcomes [20, 114, 128, 111, 139, 107, 68, 97, 6]. Examples of social determinants include environmental impact on health, educational resource in the neighboring area, economic profile of the geolocation, collective values of the community, and so on.
The core distinction between sensitive attributes and social determinants does not lie in whether an attribute is measured at the individual level, but rather in whether it is intrinsic to the individual. In other words, the distinction hinges on whether the attribute is determined solely by the individual or shaped by contextual influences from the surrounding environment. For instance, while the commute time is an individual-specific variable, it is not intrinsic to the person and depends on external factors such as transportation infrastructure. A person’s commute time can be dramatically different when this same person moves to a different neighborhood. Although a social determinant is not typically regarded as a sensitive attribute under legal or moral frameworks, it nevertheless has direct influence on one’s overall wellbeing.
2.2 Examinations of Social Determinants in Adjacent Disciplines
We provide a high-level summary of discussions on social determinants from related disciplines, including political philosophy, economics and sociology, and healthcare.
@@ 2.2 Examinations of Social Determinants in Adjacent Disciplines
In political philosophy, researchers have proposed to shift from a focus on distributive patterns to procedural issues of participation in deliberation and decision-making, and to consider structural injustices that arise from individuals’ relations to contextual environments and social institutions [12, 48, 13, 137, 138, 139]. Recent works in algorithmic fairness have urged a shift beyond the localized concerns of distributive justice, advocating for the need to investigate structural injustice [63], and to explicitly incorporate procedural inquires into fairness assessments [50, 147, 126]. However, despite growing recognition of the relevance of structural injustice and social determinants, they remain insufficiently examined in the algorithmic fairness literature. In particular, the insights from political philosophy concerning structural injustice have not been integrated to the same extent as those related to distributive justice [99, 100]. While political philosophy provides the theoretical foundation for understanding structural injustice, economics and sociology offer quantitative indices to measure these concepts.
@@
In the effort of quantitatively measuring social determinants, economists and sociologists have proposed various indices to capture the influence of contextual environments on individuals’ opportunities and outcomes. For instance, the Social Vulnerability Index (SVI) is developed by CDC/ATSDR to address social vulnerability as it relates to natural or human-caused hazards and public health emergencies [61, 44]. The (updated) Area Deprivation Index (ADI) and Neighborhood Atlas are developed by the University of Wisconsin-Madison to rank neighborhoods by socioeconomic disadvantage in a region of interest, e.g., at the state or national level [69, 68]. The Index of Concentration at the Extremes (ICE) measures spatial polarization of extreme privilege and deprivation [81, 71]. The Child Opportunity Index (COI) measures the quality of resources and conditions that matter for children’s healthy development in the neighborhoods where they live [1].
@@
Utilizing indices proposed in economics and sociology, the social determinants of health (SDoH) have long been engaged in the healthcare literature. Researchers have pointed out that one size does not fit all when it comes to the index of socioeconomic status in healthcare [15], and that the incorporation of SDoH indices necessitates methodological clarity [45]. Previous works have found racial/ethnic and geographic variations in distrust of physicians in the US [7]. The distinction between healthcare costs and healthcare needs matters when it comes to the choice of target in the development of prediction algorithms [91], and so does the distinction between race-based and race-conscious medicine [92, 26]. Most recently, the World Health Organization (WHO) has released a report about the persisting social injustices [132].
2.3 Algorithmic Fairness Approaches on Sensitive Attributes and Social Determinants
In this subsection, we review and reflect on the algorithmic fairness literature with specific attention to the treatment of sensitive attributes and social determinants.
2.3.1 Intersectionality Through Structured Combination of Sensitive Attributes
In terms of the specification of the disadvantaged individuals, previous quantitative approaches in algorithmic fairness literature primarily focus on sensitive attributes [105, 75, 35, 84, 89, 130, 21, 30, 77, 82, 145, 94, 125]. In addition to fairness notions that are applied one sensitive attribute at a time [17, 148, 62, 52, 140, 67, 72, 86, 28], intersectional fairness considerations have been introduced to account for the intersection of multiple sensitive attributes [37, 16, 66, 55, 46, 70]. While a structured combination of multiple sensitive attributes provides a more nuanced characterization of intersecting factors in discrimination, it does not fully capture the influence from contextual environments. For instance, individuals with an identical configuration of sensitive attributes, e.g., the group of African American women, can face different levels of structural injustice depending on the contextual environments they are subjected to [7, 91, 92].
2.3.2 Causal Modeling of Discrimination Originating Solely from Sensitive Attributes
In terms of the modeling of the instantiation of discriminations, previous causal fairness approaches represent discriminations with edges or pathways in the causal graph [67, 72, 86, 28, 133, 36, 38, 90, 88]. These objectionable edges or pathways typically originate from sensitive attributes [67, 72, 86, 28, 133]. However, the implications and interpretations of this technical choice may not always align with the intention to better understand and characterize the underlying causal mechanisms behind discriminations. Let us consider the usage of a directed causal edge $\texttt{Race}\rightarrow\texttt{Education Status}$ to capture racial discrimination in education [72, 86, 28, 88], as an example.
First, from the perspective of ontological and epistemological conditions, the utilization of counterfactuals can require an incoherent theory of what sensitive attributes are [64]. Here, the “utilization of counterfactuals” refers to the practice of formulating fairness notions by considering alternative values of the sensitive attribute as the basis for comparison, followed by incorporating bounds or estimations of causal effects [67, 72, 86, 143, 142, 28, 133, 59, 83, 36, 38, 90, 88]. There are different positions about what race is (e.g., the geo-biological essentialism, the racial skepticism, and the social constructionism) [53, 49, 51]. As a result, the technical ways to generate and evaluate counterfactuals involve making presumptions and choices, which require greater caution than is typically exercised in current approaches [64].
Second, from the technical perspective of causal modeling, the interpretation of the edge according to the definition of causal intervention may unintentionally recapitulate existing stereotype. Specifically, by definition of causality [116, 93], this edge asserts that there is a difference in the distribution of education status, when we “intervene” on individual’s race while keeping all other things unchanged.222 Here, we use “intervene” in quotes to signify the need of extra caution when discussing the manipulation of individual’s race, due to both ethical and practical considerations. The seemingly neutral technical treatment may unintentionally align with the controversial ideology of racial essentialism (racial groups possess underlying intrinsic essences, e.g., intellectual and biological, that make them different), which has been widely criticized due to the lack of scientific evidence supporting its claims [103, 113, 39]. The reductive summary of the instantiation of discrimination into edges or pathways originating from sensitive attributes may induce thought inertia or a force of habit. While natural and intuitive, this approach can potentially overshadow other critical contributing factors and alternative perspectives, such as the influence of social determinants.
2.3.3 Data Processing Practices and Benchmarks
@@
Other than individual-level variables, contextual environments actually have significant influences over the individual [20, 114, 128, 111, 139, 107, 68, 97, 6]. For instance, for the variable Address (or its alternatives), the improvement in physical health was observed in a randomized housing mobility social experiment [76], and the social determinants of health are closely related to individual’s residence area [79, 14, 104, 135].
@@
However, in the algorithmic fairness literature, it is a common practice to omit variables (e.g., Address) that do not directly pertain to individuals, when performing the prediction or decision-making tasks of interest. For instance, previous causal fairness approaches do not include address-related variables when modeling the data generating process with a causal graph [67, 72, 86, 143, 142, 28, 133, 59, 83, 36, 38, 90, 88]. The empirical approaches to enforce various fairness notions also tend to drop address information during data collection and/or preprocessing. Specifically, there is no address information included in the Adult dataset [11], which is widely used for evaluation purposes [17, 141, 2, 86, 41, 3, 9]. Although the Communities and Crimes dataset [101] initially contains geolocation, such information is dropped when processing the data [80]. The address information is dropped by the Folktables package when retrieving public-use data products from US Census Bureau and constructing Adult-like prediction tasks [40].
2.3.4 Remark
Algorithmic fairness analyses are naturally interdisciplinary [65, 10]. However, while the concept and importance of social determinants, in addition to sensitive attributes, have been extensively discussed in various fields (Section 2.2) and also in qualitative ways in algorithmic fairness [64, 112, 117], the quantitative approaches have primarily focused on sensitive attributes.
3 Theoretical Characterization: College Admission as a Concrete Setting
In this section, through a concrete setting of college admissions, we demonstrate how incorporating social determinants, even with geographic region as a simple surrogate, recovers nuanced insights that resonate with prior qualitative debates in jurisprudence.333 For the purpose of this paper, we aim to demonstrate how our theoretical and quantitative analyses dovetail ethical and legal insights, and we do not intend to make any legal claim. In Section 3.1, we present a summary of the assumptions we use to facilitate closed-formula theoretical analyses. In Sections 3.23.4, we consider three mainstream college admission procedures.444 We provide background information about the college admission procedures and present proofs for our theoretical results in Appendix B.

### v2

# Supplement to “Position: Beyond Sensitive Attributes, ML Fairness Should Quantify Structural Injustice via Social Determinants”

@@ Abstract
Algorithmic fairness research has largely framed unfairness as discrimination along sensitive attributes. However, this approach limits visibility into unfairness as structural injustice instantiated through social determinants, which are contextual variables that shape attributes and outcomes without pertaining to specific individuals. This position paper argues that the field should quantify structural injustice via social determinants, beyond sensitive attributes. Drawing on cross-disciplinary insights, we argue that prevailing technical paradigms fail to adequately capture unfairness as structural injustice, because contexts are potentially treated as noise to be normalized rather than signal to be audited. We further demonstrate the practical urgency of this shift through a theoretical model of college admissions, a demographic study using U.S. census data, and a high-stakes domain application regarding breast cancer screening within an integrated U.S. healthcare system. Our results indicate that mitigation strategies centered solely on sensitive attributes can introduce new forms of structural injustice. We contend that auditing structural injustice through social determinants must precede mitigation, and call for new technical developments that move beyond sensitive-attribute-centered notions of fairness as non-discrimination.
Machine Learning, ICML
@@ 1 Introduction
Structural injustice refers to circumstances in which social practices, social structures, or the environment reinforce and compound prior histories of injustice (Carmichael et al., 1967; Sowell, 1972; Young, 1990; Tilly, 1998; Rothstein, 2017; Alexander, 2020). We use the term “social determinants” to denote the specific, measurable aspects of these practices, structures, or environments (i.e., the variables or features) through which structural injustice manifests and can be systematically audited (Gee & Ford, 2011; Yearby, 2018; Robinson et al., 2020; Yearby et al., 2022; Chetty et al., 2024). In contrast to auditing unfairness as structural injustice instantiated through social determinants, the machine learning (ML) fairness literature has largely operationalized auditing unfairness as discrimination along sensitive attributes, e.g., race, sex, gender, and age (Romei & Ruggieri, 2014; Loftus et al., 2018; Corbett-Davies & Goel, 2018; Mitchell et al., 2018; Narayanan, 2018; Verma & Rubin, 2018; Caton & Haas, 2020; Chouldechova & Roth, 2020; Makhlouf et al., 2020; Mehrabi et al., 2021; Zhang & Liu, 2021; Pessach & Shmueli, 2022; Tang et al., 2023). In this position paper, we argue that beyond sensitive attributes, ML fairness should quantify structural injustice via social determinants.
@@
Because social determinants are features of places, institutions, policies, or practices, unfairness as structural injustice persists even if animuses that cause unjust treatment (unfairness as discrimination) have been subject to significant reform. Their effects may not be tied directly to demographic group membership but to broader traits (such as income level or job type) or to geographic areas. As a result, individuals within the same demographic group, depending on their unique circumstances, may experience different levels of (dis)advantage due to intersecting social determinants, e.g., various environmental impacts on health in terms of distance and accessibility to general practitioners and hospitals (Comber et al., 2011), Vitamin D and bone health across geographic locations (Yeum et al., 2016), and geographic disparities of diseases (Tan et al., 2020). Conversely, individuals from different demographic groups in the same geographic neighborhood may encounter similar impediments, e.g., poverty and pollution in the neighborhood, lack of educational resource in the community (Connell, 1994; Tilak, 2002; Rose & Dyer, 2008).
Auditing constitutes the essential first step in shaping the fairness problem space within which mitigation strategies operate. To support our position, our argument unfolds as follows. (I) Conceptual Gap: We identify the gap between cross-disciplinary treatments of social determinants for structural injustice, and their limited engagement in ML fairness research (Section 2). (II) Paradigm Limits: We show that the prevailing sensitive-attribute-centered technical paradigms are not well suited to quantifying structural injustice via social determinants (Section 3). (III) Unintended Injustice: We theoretically demonstrate that mitigation strategies focused on sensitive attributes can introduce additional structural injustice (Section 4). (IV) Empirical Evidence: We provide empirical evidence for the necessity and value of social-determinant-based analysis and intervention (Section 5). (V) Actionable Reorientation: We engage alternative perspectives (Section 6) and propose concrete actions for addressing the limitations of the current paradigm (Section 7).
@@ 2 Cross-Disciplinary Engagement with Social Determinants
In this section, we review and reflect on the cross-disciplinary engagement with social determinants. [^1] Due to space limit, we provide further discussions on related works in Appendix A. In Section 2.1, we clarify definitions of sensitive attributes and social determinants. In Section 2.2, we summarize discussions on social determinants across disciplines.
2.1 Define Sensitive Attributes and Social Determinants
@@
A social determinant is a variable representing an aspect of the data generating process, in which one or more characteristics of the context (e.g., social practices/structures, environments) where individuals live or operate, that are not an attribute of any specific individual, have direct influence on their attributes.
Sensitive attributes are relatively stable identifiers arising from longstanding legal and moral frameworks, that can be uniquely ascribed to an individual, e.g., race, sex, disability status. Social determinants refer to external conditions that influence an individual’s opportunities, behaviors, and outcomes (Carmichael et al., 1967; Sowell, 1972; Tilly, 1998; Singh, 2003; Young, 2008; Rothstein, 2017; Kind & Buckingham, 2018; Powers & Faden, 2019; Alexander, 2020; Kasirzadeh & Smart, 2021; Smart & Kasirzadeh, 2024; Sullivan & Kasirzadeh, 2024). Examples of social determinants include environmental impact on health, educational resource in the neighboring area, economic profile of the geolocation, collective values of the community, and so on.
2.2 Engagement Across Disciplines
In this subsection, we provide a high-level summary of discussions on social determinants from related disciplines of algorithmic fairness, including political philosophy, economics and sociology, and healthcare.
@@ 2.2 Engagement Across Disciplines
In political philosophy, researchers have proposed to shift from a focus on distributive patterns to procedural issues of participation in deliberation and decision-making, and to consider structural injustices that arise from individuals’ relations to contextual environments and social institutions (Blau, 1977; Giddens, 1979; Bourdieu, 1984; Young, 1990, 2006, 2008). Recent works in algorithmic fairness have urged a shift beyond the localized concerns of distributive justice, advocating for the need to investigate structural injustice (Kasirzadeh, 2022), and to explicitly incorporate procedural inquires into fairness assessments (Grgić-Hlača et al., 2018; Zimmermann & Lee-Stronach, 2022; Tang et al., 2024).
@@
In the effort of quantitatively measuring social determinants, economists and sociologists have proposed various indices to capture the influence of contextual environments on individuals’ opportunities and outcomes. For instance, the (updated) Area Deprivation Index (ADI) and Neighborhood Atlas are developed to rank neighborhoods by socio-economic disadvantage in a region of interest, e.g., at the state or national level (Kind et al., 2014; Kind & Buckingham, 2018). The Index of Concentration at the Extremes (ICE) measures spatial polarization of extreme privilege and deprivation (Massey, 2001; Kubrin & Stewart, 2006). The Child Opportunity Index (COI) measures the quality of resources and conditions that matter for children’s healthy development in the neighborhoods where they live (Acevedo-Garcia et al., 2014).
@@
The social drivers of health (SDoH) have long been engaged in the healthcare literature. Researchers have pointed out that one size does not fit all when it comes to the index of socio-economic status in healthcare (Braveman et al., 2005), and that the incorporation of SDoH indices necessitates methodological clarity (Foryciarz et al., 2025). Previous works have found racial/ethnic and geographic variations in distrust of physicians in the U.S. (Armstrong et al., 2007). The distinction between healthcare costs and healthcare needs matters when it comes to the choice of target in the development of prediction algorithms (Obermeyer et al., 2019), and so does the distinction between race-based and race-conscious medicine (Pallok et al., 2019; Cerdeña et al., 2020). Most recently, the World Health Organization (WHO) has released a report about the persisting social injustices (World Health Organization, 2025).
Takeaway: Despite extensive cross-disciplinary work on structural injustice instantiated through social determinants, treatment within ML fairness remains limited.
3 Limitations of Current Technical Paradigms
In this section, we argue that current technical paradigms are inadequate for capturing unfairness as structural injustice.
3.1 Existing Practices and Benchmarks Tend to Drop Attributes Related to Social Determinants
@@
In addition to individual-level variables, contextual environments have significant influences over the individual (Carmichael et al., 1967; Sowell, 1972; Tilly, 1998; Singh, 2003; Young, 2008; Rothstein, 2017; Kind & Buckingham, 2018; Powers & Faden, 2019; Alexander, 2020; Perdomo et al., 2025). For instance, for the variable Address (or its alternatives), the improvement in physical health was observed in a randomized housing mobility social experiment (Ludwig et al., 2011), and the social determinants of health are closely related to individual’s residence area (Marmot & Wilkinson, 2005; Braveman & Gottlieb, 2014; Robinson et al., 2020; Yearby et al., 2022).
@@
However, in ML fairness literature, it is a common practice to omit variables that do not directly pertain to individuals, when performing the prediction or decision-making tasks of interest. For instance, previous causal fairness approaches do not include address-related variables when modeling the data generating process with a causal graph (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Zhang & Bareinboim, 2018b, a; Chiappa, 2019; Wu et al., 2019; Imai & Jiang, 2020; Mishler et al., 2021; Coston et al., 2020; D’Amour et al., 2020; Nilforoshan et al., 2022; Nabi et al., 2022). Although the Communities and Crimes dataset (Redmond, 2009) initially contains geolocation, such information is dropped during data processing (Mary et al., 2019).
Moreover, widely used benchmark datasets typically exclude variables that potentially capture social determinants. For instance, there is no address information included in the Adult dataset (Becker & Kohavi, 1996), which is a standard dataset for evaluation purposes (Calders et al., 2009; Zemel et al., 2013; Agarwal et al., 2018; Nabi & Shpitser, 2018; Donini et al., 2018; Agarwal et al., 2019; Baharlouei et al., 2020). Similar to the attribute list of Adult dataset, the address information is dropped by the Folktables package when retrieving public-use U.S. census data products to construct Adult-like prediction tasks (Ding et al., 2021), except for the ACSTravelTime task.
3.2 (Quasi-)Stable Sensitive Attributes Cannot Capture Shifting Influences from Social Determinants
In terms of the specification of the disadvantaged individuals, previous quantitative approaches in algorithmic fairness literature primarily focus on sensitive attributes. In addition to fairness notions that are applied one sensitive attribute at a time (Calders et al., 2009; Žliobaite et al., 2011; Kamiran et al., 2013; Hardt et al., 2016; Zafar et al., 2017; Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019), intersectional fairness considerations have been introduced to account for the intersection of multiple sensitive attributes through their structural combinations (Crenshaw, 1990; Bright et al., 2016; Kearns et al., 2018; Hoffmann, 2019; Foulds et al., 2020; Kong, 2022).
While a structured combination of multiple (quasi-)stable sensitive attributes provides a more nuanced characterization of intersecting factors in discrimination, it does not fully capture the shifting influence from contextual environments, in static and and dynamic settings. For instance, individuals with an identical configuration of sensitive attributes, e.g., the group of African American women, can face different levels of structural injustice depending on the contextual environments they are subjected to (Armstrong et al., 2007; Obermeyer et al., 2019; Pallok et al., 2019).
These influences are not intrinsic to the individual. Instead, they characterize the surrounding environment and dynamically shape the conditions under which outcomes are produced. Consequently, beyond asking which attributes explain (residual) dependence between individual’s relatively stable sensitive attributes and outcomes (e.g., Conditional Demographic Parity by Žliobaite et al. 2011; Kamiran et al. 2013 and further discussed by Wachter et al. 2021), fairness metrics should also adaptively and dynamically account for benefits/burdens induced by contextual variation such as neighborhood-level differences in healthcare access, exposure, and institutional resources.
3.3 Mismatch Between Individual-Level Causal Modeling and Community-Level Structure
In terms of the modeling of the instantiation of unfairness, previous causal fairness approaches represent discriminations with edges or pathways in the causal graph (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019; Wu et al., 2019; Coston et al., 2020; D’Amour et al., 2020; Nilforoshan et al., 2022; Nabi et al., 2022; Zuo et al., 2022), which typically originate from sensitive attributes of an individual. Social determinants, especially community- or context-level attributes (e.g., neighborhood deprivation or healthcare access), create a fundamental mismatch with most causal modeling frameworks in ML fairness, which primarily operate at the individual level.
Specifically, individual-level graph formulations do not readily represent community-level variables whose values emerge from shared environments and population composition. Many social determinants are aggregate statistics over populations, inducing potentially cyclic causal influence between individuals and communities. Individual-level causal interventions can propagate to the community level, and community conditions in turn shape individual outcomes. For example, intervening on an individual’s income can shift neighborhood socio-economic status, affecting others in the community and violating standard assumptions such as no interference (Hernán & Robins, 2020) that underlie much of causal inference and fairness analysis.
Existing causal fairness notions further reinforce this individual-level structure by analyzing path-specific causal effects within individual-level graphs. Capturing effects of social determinants requires modeling additional structure across individuals and communities, which standard individual-level models do not encode. Although some work aggregates individual causal effects across subgroups (Coston et al., 2020; Imai & Jiang, 2020; Mishler et al., 2021), the underlying models remain individual-level, without capturing the bidirectional influences between an individual and their communities. Consequently, current causal fairness approaches are ill-suited to represent unfairness instantiated through collective, community-level social determinants.
Takeaway: Existing technical paradigms, exemplified by standard benchmarks, sensitive-attribute-centered fairness metrics, and causal models of unfairness as discrimination, are inadequate for capturing structural injustice.
4 Tension Between Structural Justice and Sensitive-Attribute-Centered Mitigation
To formalize the intuition that identity-based metrics fail to capture structural context, we introduce a stylized model and analyze college admissions as an instance of sensitive-attribute-centered mitigation grounded in U.S. affirmative-action efforts to remedy race-based historical injustice. We theoretically demonstrate that incorporating social determinants, even through a simple geographic proxy, not only recovers nuances consistent with prior qualitative discussions, but also uncovers a counter-intuitive paradox between sensitive-attribute-centered mitigation and progress in structural justice. In Section 4.1, we present assumptions we use to facilitate closed-formula theoretical analyses. In Sections 4.2, we demonstrate structural injustice implications of a classic sensitive-attribute-centered mitigation strategy. [^2] Additional results and proofs are provided in Appendix B.